The activity of a sample of radioactive material is 'X' at time t and 'Y' at time 3t,If it's mean life is 'T' ,then.......


1) X = 3Y


2) X - Y = 2t


3) Y = X e^(-2t/T)


4) Y = X e^(-3t/T)


 


Pls explain......

Grade:12

1 Answers

aman gupta
13 years ago

mean life = 1/λ=T

activity= λn

i) X= λNe-λt

ii) Y=λNe-λ3t

divide (i)and (ii)

ans 3